Determine the concentrators of Na_2CO_3 Na^+. and General chemistry 2. Help please.
Answer #1

Answer – We are given, mass of Na2CO3 = 2.06*10-4 g

Volume = 2.25 L

First we need to calculate the moles of Na2CO3

We know,

Moles of Na2CO3 = 2.06*10-4 g / 105.98 g.mol-1

                              = 1.94*10-6 moles

So, 1 moles of Na2CO3 =1 moles of CO32-

So, 1.94*10-6 moles of Na2CO3 = ? moles of CO32-

= 1.94*10-6 moles of CO32-

1 moles of Na2CO3 = 2 moles of Na+

So, 1.94*10-6 moles of Na2CO3 = ?

= 3.88*10-6 moles of Na+

So, molarity of Na2CO3 = 1.94*10-6 moles / 2.25 L

                                      = 8.64*10-7 M

Molarity of Na+ = 3.88*10-6 moles / 2.25 L

                          = 1.73*10-6 M

molarity of CO32- = 1.94*10-6 moles / 2.25 L

                           = 8.64*10-7 M

We know, 1 ppm = 1 mg/ L

So, mass of Na+ = 3.88*10-6 moles * 22.990 g/mol

                           = 8.93*10-5 g

                           = 0.0893 mg

So, pppm of Na+ = 0.0893 mg / 2.25 L

                            = 0.0397 ppm

Mass of CO32- = 1.94*10-6 moles * 60.08 g/mol

              = 0.000117 g

                = 0.117 mg

So, ppm of CO32- = 0.117 mg / 2.25 L

                            = 0.0519 ppm